# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v11

@@ 1 Introduction
This study’s primary contribution is to add analytical perspectives on consumption smoothing behavior to economic history studies. Studies of how working-class households compensated for income loss during the late 19th to the early 20th century have focused on Britain. These studies used cross-sectional variation in income to estimate the elasticities of the temporary income sources. Their detailed analyses have revealed the roles of financial institutions and labor supply adjustments in compensating for income loss in working-class households (Scott and Walker 2012). [^2] They found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use savings institutions. Horrell and Oxley (2000) highlight the important role of labor supply adjustments among family members to manage the income losses of the head in working-class households in late 19th-century Britain. Kiesling (1996) and Boyer (1997) focused on the risk-mitigation strategies among the cotton workers households in Lancashire during the recession caused by the Civil War. However, there is little evidence of how those risk-coping strategies smoothed their consumption. [^3] Other strands of literature provide supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005), building image on the historical consumers in the long-run (Francks 2009), and describe the relationship between consumption and production (Gordon 2013). Hence, this study provides a rare window into urban household financial behavior during industrialization by utilizing unique household longitudinal budget data from prewar Japan. It also implements a systematic panel data analysis to offer comparable estimates of the income elasticity for future studies.
@@ 5 Credit System in Local Economy
The foregoing results enlighten the roles of credit institutions among skilled workers’ households during industrialization. As the credit for purchases was provided by local retailers, this suggests the importance of the relationships between the workers and retailers. To the best of my knowledge, however, there is little historical research on credit purchases in prewar Japan. Hence, the details of this credit institution are not yet clear.
Table 9: Customers, Sales Methods, and Competitiveness in the Retailers in Tokyo City
White Rice 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
Fish 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
Greengrocer 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
Firewood and charcoal 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
The information on the credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented. Fortunately, however, I found that some official survey reports documented information on this credit institution. A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers. [^69] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p. 1). Those included retailers as well as wholesale and other small businesses. The number of respondents varied depending on the question items. The number of respondents in the statistical table on the payment methods was $2,654$, of which $1,197$ were retailers (Tokyo City Office 1932b, p. 273). To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo circa 1930. In fact, $99$%($155/157$) of grain retailers, $81$%($35/43$) of fish retailers, $78$%($64/82$) of greengrocers, and $94$%($45/48$) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp. 260; 263). Thus, credit purchases in Tokyo were neither region- nor religion-specific customs. A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information. Table 9 summarizes the information on the retailers’ sales measured in this survey. [^70] This survey covered $939$ retailers, and its report was edited on December 31, 1935. I digitized the statistics for all the retailers selling the representative items (rice, fish, fruit and vegetables, and fuels). The survey instrument included detailed questions regarding sales and purchases (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by scale of businesses. I confirm that the main findings are materially similar. First, many annual sales were from their regular customers, most of whom were from the non-upper classes. For example, $73$% of annual sales of rice retailers were from their regular customers, and more than $80$% of them were from non-upper classes, including factory workers. Second, credit purchase was a popular sales method, especially among rice and fuel retailers—more than $70$% of the sales were conducted with credit in both shops. [^71] These figures are systematically lower than those from the 1930 Survey Report. Given that both surveys show a systematic tendency toward credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself. Although cash was also used among greengrocers, $34$% of the sales were settled using credit. This suggests that credit purchase was a popular payment method and that regular customers settled a large part of it. [^72] The 1930 Survey reveals that the retailers also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp. 154–157). This means that credit purchase was a familiar payment method for retailers as well as consumers. Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions with long-term relationships with their regular customers. [^73] A textbook for retailers published in 1924 explains the psychology behind this long-term relationship—credit purchase was provided for the convenience of the customer’s budget but also for the “vanity” (status) of being regular customers (Shimizu 1924, pp. 115–116). The final column of this table further provides suggestive evidence for this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (approximately $330$ meters) radius. As shown, rice and fuel stores have two to three more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
@@
Next, I examined the credit purchases included in the THBS dataset. Most THBS households ($29/33$) used any credit purchases during their sample period, confirming that credit purchase was a common practice among factory workers. Figure 6a illustrates the average daily credit purchases and payments. Generally, there is a peak per semi-month in payments, which reflects the practice of making the payments for the credit purchase on or soon after their paydays. The purchase series shows a similar tendency but with lower spikes and larger variance. This indicates that the households first purchased goods for a certain amount on or soon after their paydays, and then used purchased perishable foods such as fish, fruit, and vegetables when they needed them. Figure 6b illustrating the net purchase confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays. Figure 6c shows that the payment slightly exceeded purchase, suggesting that many households could approximately repay their loans. Although the difference between payment and purchase is relatively obvious after August 1919, Figure 6d indicates that this difference is not statistically significant in all the semi-months. [^74] Strictly, payment can include accrued liabilities that should have been paid prior to the previous period. It is, therefore, reasonable that the average payment is greater than the average purchase in most semi-month cells. Thus, this is not a systematic trend but may be due to the smaller available number of households in the later sample period (Online Appendix Figure B.4).
@@
$I(\overline{CP}_{i}=0)=\varkappa_{0}+\varkappa_{1}CV_{i}+\bar{\mathbf{x}}_{i}^{\prime}\mathbf{\chi}+v_{i},$ (4)
@@
This, in turn, implies that retailers may factor risk into retail prices or by charging interest on their loans. Unfortunately, there are no systematic statistics on the interest charged on credit purchases. It may be possible to reverse calculate the risk premiums by differencing payments and purchases. This reverse calculation technique requires information on both the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were frequently sold in greengrocers (yaoya) but sometimes also in general stores (aramonoya). Fortunately, I found a specific month in a budget book that has documented the price-by-retailer information. [^76] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and general stores between February 1 and February 14, 1919, and paid exactly the same amounts on the 15th. [^77] The repayment period in credit purchases among retailers and wholesalers (footnote 72) was often set at approximately one month (Tokyo City Office 1932, pp. 157-159). Therefore, customers likely made repayments at semi-monthly or monthly intervals. In the latter half of February, the household paid the same amounts to greengrocers and general stores on the 28th, but paid less to the fish stores. As the budget book for March does not have information on the type of retailers, it is difficult to determine exactly how the reimbursement shortfall was managed. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
@@
Finally, I assessed the losses suffered by retailers for credit purchases. According to the 1930 Survey Report, almost all the retailers suffered losses for credit purchases. For example, $99$% of rice and $97$% of fuel retailers lost their sales. This means that credit purchase was inherently a loss-making sales method for the retailers. As systematic statistics are unavailable to clarify the percentage of total profits corresponding to these losses, it is impossible to assess how severe these losses were for the retailers. However, an important fact is that these losses were usually capped at 5% of total sales. For example, the available statistics suggest that $87$% of rice and $65$% of fuel retailers suffered losses amounting to $5$% of sales. This supports the evidence that while the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent. This seems to be a plausible sales strategy for retailers to manage adverse selection issues under imperfect consumer information. Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail.
@@ 6 Conclusion
This study investigates the consumption-smoothing strategies used by working-class households in prewar Tokyo. I digitized the daily longitudinal budget survey on factory worker households conducted in Tsukishima in 1919. I carefully assessed the potential measurement error due to the miss-assignments of paydays and proposed a way to avoid the assignment errors. The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks.
The estimated elasticity for total expenditure ($0.43$) is in a similar range to those in rural economies in developing countries. According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was approximately $0.5$. As for the urban households, the estimated elasticity was $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). Since my estimate is from a specific occupation, skilled factory workers, we need to be careful when making comparisons with these estimates based on urban samples which include a variety of occupations. Nevertheless, the factory workers in the historical metropolis of Japan circa WWI were vulnerable to idiosyncratic income shocks. A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024). My estimate from the adjusted monthly panels is slightly larger than this estimate. If taken literally, this implies that the Tsukishima factory workers were slightly more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind consumption are subjects for future work.
@@ B.1 Tsukishima Survey: The First Social Survey in Japan
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916868686This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p. 33). The original documents of this survey, preserved at Tokyo Imperial University, were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). Before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.878787The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason was the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the seventh subcommittee evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp. 55–59). [^88] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda (2015, pp. 76–77) used the diary to analyze the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network. On October 26, Tanahashi and Yamana—together with the Kyobashi Branch of the fraternity club—conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection there on October 30 and decided on Tsukishima as the survey site.

### v12

@@ 1 Introduction
This study’s primary contribution is to add analytical perspectives on consumption smoothing behavior to economic history studies. Studies of how working-class households compensated for income loss during the late 19th to the early 20th century have focused on Britain. These studies used cross-sectional variation in income to estimate the elasticities of the temporary income sources. Their detailed analyses have revealed the roles of financial institutions and labor supply adjustments in compensating for income loss in working-class households (e.g., Scott and Walker 2012). [^2] They found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use savings institutions. Horrell and Oxley (2000) highlight the important role of labor supply adjustments among family members to manage the income losses of the head in working-class households in late 19th-century Britain. Kiesling (1996) and Boyer (1997) focused on the risk-mitigation strategies among the cotton workers households in Lancashire during the recession caused by the Civil War. However, there is little evidence of how those risk-coping strategies smoothed their consumption. [^3] Other strands of literature provide supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005), building image on the historical consumers in the long-run (Francks 2009), and describe the relationship between consumption and production (Gordon 2013). Hence, this study provides a rare window into urban household financial behavior during industrialization by utilizing unique household longitudinal budget data from prewar Japan. It also implements a systematic panel data analysis to offer comparable estimates of the income elasticity for future studies.
@@ 5 Credit System in Local Economy
The foregoing results enlighten the roles of credit institutions among skilled workers’ households during industrialization. As the credit for purchases was provided by local retailers, this suggests the importance of the relationships between the workers and retailers. To the best of my knowledge, however, there is little historical research on credit purchases in prewar Japan. Hence, the details of this credit institution are not yet clear. This may mainly be because the information on credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented. Accordingly, I analyze the institution using both the THBS dataset and available official documents.
5.1 Availability and Premiums: Evidence from the THBS Dataset
@@
I begin this descriptive analysis by examining the credit purchases included in the THBS dataset. Most THBS households ($29/33$) used any credit purchases during their sample period, confirming that credit purchase was a common practice among factory workers. Figure 6a illustrates the average daily credit purchases and payments. Generally, there is a peak per semi-month in payments, which reflects the practice of making the payments for the credit purchase on or soon after their paydays. The purchase series shows a similar tendency but with lower spikes and larger variance. This indicates that the households first purchased goods for a certain amount on or soon after their paydays, and then used purchased perishable foods such as fish, fruit, and vegetables when they needed them. Figure 6b illustrating the net purchase confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays. Figure 6c shows that the payment slightly exceeded purchase, suggesting that many households could approximately repay their loans. Although the difference between payment and purchase is relatively obvious after August 1919, Figure 6d indicates that this difference is not statistically significant in all the semi-months. [^69] Strictly, payment can include accrued liabilities that should have been paid prior to the previous period. It is, therefore, reasonable that the average payment is greater than the average purchase in most semi-month cells. Note again that the wider variances in the latter sample period are based on the smaller available number of households (Online Appendix Figure B.4).
@@
$I(\overline{CP}_{i}=0)=\varkappa_{0}+\varkappa_{1}CV_{i}+\bar{\mathbf{x}}_{i}^{\prime}\bm{\chi}+v_{i},$ (4)
@@
This, in turn, implies that retailers may factor risk into retail prices or by charging interest on their loans. Unfortunately, there are no systematic statistics on the interest charged on credit purchases. It may be possible to reverse calculate the risk premiums by differencing payments and purchases. This reverse calculation technique requires information on both the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were frequently sold in greengrocers (yaoya) but sometimes also in general stores (aramonoya). Fortunately, I found a specific month in a budget book that has documented the price-by-retailer information. [^71] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and general stores between February 1 and February 14, 1919, and paid exactly the same amounts on the 15th. [^72] The repayment period in credit purchases among retailers and wholesalers (footnote 76) was often set at approximately one month (Tokyo City Office 1932, pp. 157-159). Therefore, customers likely made repayments at semi-monthly or monthly intervals. In the latter half of February, the household paid the same amounts to greengrocers and general stores on the 28th, but paid less to the fish stores. As the budget book for March does not have information on the type of retailers, it is difficult to determine exactly how the reimbursement shortfall was managed. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
5.2 Local Environment and Retailers’ Incentives: Evidence from the Official Surveys
Next, I assess circumstances among the retailers using available survey reports on the credit institution. A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers. [^73] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p. 1). Those included retailers as well as wholesale and other small businesses. The number of respondents varied depending on the question items. The number of respondents in the statistical table on the payment methods was $2,654$, of which $1,197$ were retailers (Tokyo City Office 1932b, p. 273). To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo circa 1930. In fact, $99$%($155/157$) of grain retailers, $81$%($35/43$) of fish retailers, $78$%($64/82$) of greengrocers, and $94$%($45/48$) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp. 260; 263). Thus, credit purchases in Tokyo were neither region- nor religion-specific customs.
Table 9: Customers, Sales Methods, and Competitiveness
in the Retailers in Tokyo City
White Rice 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
Fish 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
Greengrocer 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
Firewood and charcoal 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information. Table 9 summarizes the information on the retailers’ sales measured in this survey. [^74] This survey covered $939$ retailers, and its report was edited on December 31, 1935. I digitized the statistics for all the retailers selling the representative items (rice, fish, fruit and vegetables, and fuels). The survey instrument included detailed questions regarding sales and purchases (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by scale of businesses. I confirm that the main findings are materially similar. First, many annual sales were from their regular customers, most of whom were from the non-upper classes. $73$% of annual sales of rice retailers were from their regular customers, and more than $80$% of them were from non-upper classes, including factory workers. Second, credit purchase was a popular sales method, especially among rice and fuel retailers—more than $70$% of the sales were conducted with credit in both shops. [^75] These figures are systematically lower than those from the 1930 Survey Report. Given that both surveys show a systematic tendency toward credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself. Although cash was also used among greengrocers, $34$% of the sales were settled using credit. This suggests that credit purchase was a popular payment method and that regular customers settled a large part of it. [^76] The 1930 Survey reveals that the retailers also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp. 154–157). This means that credit purchase was a familiar payment method for retailers as well as consumers. This confirms that the retailers were engaged in credit transactions with long-term relationships with their regular customers. [^77] A textbook for retailers published in 1924 explains the psychology behind this long-term relationship—credit purchase was provided for the convenience of the customer’s budget but also for the “vanity” (status) of being regular customers (Shimizu 1924, pp. 115–116). The final column of this table further provides suggestive evidence for this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (approximately $330$ meters) radius. As shown, rice and fuel stores have two to three more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit. Online Appendix C.3 provides a set of regression-based evidence on the positive relationship between the use of credit purchases and the number of local competitors.
@@ 5.2 Local Environment and Retailers’ Incentives: Evidence from the Official Surveys
Finally, I assessed the losses suffered by retailers for credit purchases. According to the 1930 Survey Report, $99$% of rice and $97$% of fuel retailers lost their sales. [^78] Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail. This implies that credit purchase was inherently a loss-making sales method for the retailers. However, these losses were usually marginal and capped at 5% of the total sales. For example, the available statistics suggest that $87$% of rice and $65$% of fuel retailers suffered losses amounting to $5$% of sales. While the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent. This seems to be a plausible sales strategy for retailers to manage adverse selection issues under imperfect consumer information.
@@ 6 Conclusion
This study investigates the consumption-smoothing strategies used by working-class households in prewar Tokyo. I digitized the daily longitudinal budget survey on factory worker households conducted in Tsukishima in 1919. I assessed the potential measurement error due to the miss-assignments of paydays and proposed a way to avoid the assignment errors. The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks.
The estimated elasticity for total expenditure ($0.43$) is in a similar range to those in rural economies in developing countries. The elasticity for the rural households in Indian villages between 1975 and 1984 was reported to be approximately $0.5$. As for the urban households, the estimated elasticity was $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan. Since my estimate is from a specific occupation, skilled factory workers, we need to be careful when making comparisons with these estimates based on urban samples which include a variety of occupations. Nevertheless, the factory workers in the historical metropolis of Japan circa WWI were vulnerable to idiosyncratic income shocks. A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920.808080The estimates for India, Bangkok, and Osaka are obtained from Ravallion and Chaudhuri (1997), Townsend (1995), and Ogasawara (2024), respectively. My estimate from the adjusted monthly panels is slightly larger than this estimate. If taken literally, this implies that the Tsukishima factory workers were slightly more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind consumption are subjects for future work. This would add a new analytical perspective to the study of economic history.
@@ B.1 Tsukishima Survey: The First Social Survey in Japan
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916888888This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p. 33). The original documents of this survey, preserved at Tokyo Imperial University, were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). Before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.898989The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason was the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the seventh subcommittee evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp. 55–59). [^90] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda (2015, pp. 76–77) used the diary to analyze the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network. On October 26, Tanahashi and Yamana—together with the Kyobashi Branch of the fraternity club—conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection there on October 30 and decided on Tsukishima as the survey site.